# A Comunication Framework for Compositional Generation

### v1

@@ B.1 Definitions of Holistic and Compositional encodings, and role of the NHE condition
As indicated in §3.4, NHE condition on the sender-receiver protocol just requires that the protocol be not holistic, and thus it is fully specified via definitions B.1 and B.2. Furthermore, Definition B.2 fully specifies in which way the map $\mathcal{G}\to\mathcal{M}$ successfully approximates an isomorphic reconstruction of $\mathcal{G}$, if it satisfies conditions NHE and UE.
B.2 Encodings optimizing for only UE and EE have Kolmogorov Complexity much higher than compositional ones
@@
As a manageable measure of complexity we will use Kolmogorov Complexity (KC), and we verify that in our setting, encodings satisfying UE and EE will necessarily have higher KC than compositional ones. In particular, the receiver will have more trouble reconstructing from such encodings, justifying our push to enforce the NHE condition.
Encodings of $\mathcal{D}_{train}$ satisfying UE and EE include all injective maps
@@
On the other hand, to specify a compositional encoding $\mathsf{comp}$ as in Definition B.2 requires only assigning a string of length $C_{i}$ in alphabet $V$ for every generating factor $G_{i}$, where we can take $C_{i}=\lceil\log_{|V|}N_{i}\rceil$. Thus such string requires $\approx\log_{2}N_{i}$ bits, and the whole encoding has KC given by
@@
Proposition B.3.
Under the same hypotheses on $\mathcal{G}$ as in the previous section, consider the limit $|\mathcal{D}|\to\infty$ and assume that $p=|\mathcal{D}_{train}|/|\mathcal{D}|$ stays bounded away from zero in the limit. Then the expected Kolmogorov complexity of encodings satisfying UE and EE becomes overwhelmingly larger than the one of compositional encodings:
@@
B.3 Encodings that optimize efficiency of encoding EE under UE over $\mathcal{D}_{train}$ are holistic
Example 3.1 from §3.4 may be sufficient to convince the reader intuitively of this phenomenon, however here we develop the idea of that example in higher formal detail.
@@ B.3 Encodings that optimize efficiency of encoding EE under UE over $\mathcal{D}_{train}$ are holistic
  • The requirement of ”UE over $\mathcal{D}_{train}$” means that the encodings are injective over $\mathcal{D}_{train}$.
  • The requirement of ”EE over $\mathcal{D}_{train}$” means that the encoding maps $\mathcal{D}_{train}$ into $V^{C_{\star}}$ for the minimum value of $C_{\star}$.
@@
For compositional encodings of $\mathbf{G}=[G_{1},\dots,G_{n}]$ the encoding will require to set apart separate message substrings of length $C_{i}=\lceil\log_{|V|}N_{i}\rceil$ for factor $G_{i}$ (see Def. B.2), thus the required length for the encoding will be
@@
We now discuss when the inequality signs in (14) are sharp or not, since sharp inequality $C_{0}>C_{\star}$ implies our claim that optimizing for EE under UE implies non-compositional (i.e., holistic) encodings:
@@
Proposition B.4.
Under the same running assumptions over $\mathcal{G}$, $p$, we have that optimizing only for efficiency of encodings EE under injectivity constraints UE over $\mathcal{D}_{train}$ is guaranteed to produce holistic (i.e., non-compositional) encodings if one or both of the conditions (15) and (16) holds.
B.4 Possible choices of encoded message regularization
As seen above, if we just optimize for EE and UE then we are likely to get holistic encodings (Prop. B.4), which will then make it hard for the receiver to generalize the decoding strategy outside the training set due to having overwhelmingly higher expected complexity (Prop. B.3).
This proves that it is important to incentivize condition NHE, i.e., to push encoding strategies away from overly efficient holistic encodings. In this section we discuss a few approaches to do this in practice, explaining our choice of NHE regularization.
@@ B.4 Possible choices of encoded message regularization
    Entropy maximization. Introduced by (Williams & Peng, 1991) for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
@@
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in (Ren et al., 2020).
@@
    Importantly, in our setting, to avoid high variance in estimation due to the large message space (see §C), we approximate the probability distribution of the RNN-produced message token at position $t\leq C$, denoted here $p(m_{t}|x,m_{[t-1]})$, as a $\delta$ distribution over the vocabulary $V$ by using the Gumbel-softmax trick:
@@
    In particular, it becomes clear that optimizing this $h$ becomes unfeasable with gradient methods.
    Adaptation to our setting: a more promising strategy. Our model can be seen as a different approximation of the entropy of the message. Instead of using the policy weights as probability values, we use the law of large numbers to estimate the probability of each message. For a batch $\mathcal{B}\subset\mathcal{X}$ consisting of $N_{batch}$ training examples and for given sender encoding protocol $S_{\phi}$:
@@
    Yet again, due to the exponential size of the message space, the entropy taken by this probability is still not practical. As an approximation, we will asume independence between symbol positions, allowing to get the stronger approximation:
@@
    $H(Pr(m\mid x,\phi))$ ${{=}}$ $-\sum_{t=1}^{C}\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi))\ln(Pr(m_{t}=v\mid x,\phi))$ (28)
    $\geq$ $\sum_{t=1}^{C}\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi)-Pr(m_{t}=v\mid x,\phi)^{2})\quad\text{(because $-x\ln x\geq x-x^{2}$ for $x\leq 1$)}$
    $\stackrel{{\scriptstyle(\star)}}{{=}}$ $\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)-\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi)^{2})\right)$
    $=$ $\sum_{t=1}^{C}\sum_{v\neq v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)=\sum_{t=1}^{C}Pr(m_{t}\neq m_{t}^{\prime})=\mathbb{E}[d_{H}(m,m^{\prime})]$
    (${{\approx}}$ $\mathbb{E}_{x,x^{\prime}\in\mathcal{B}}[d_{H}(S_{\phi}(x),S_{\phi}(x^{\prime})],\quad\quad\text{whith $\mathbb{E}$ taken w.r.t. the uniform measure over }\mathcal{B}.$
@@
    $\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)=1=\left(\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\right)^{2}=\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi).$
    The consequence of bound (28) is that maximizing the pairwise Hamming distance between predicted messages in a batch also maximizes a lower bound to the approximate entropy of the message space, justifying the choice of optimizing this quantity (which is easier to i include in practice) for favoring NHE.
@@
    Contrastive Learning losses. Contrastive loss functions such as the triplet loss (Schroff et al., 2015) and NT-Xtent (Chen et al., 2020) work by defining positive and negative image pairs (usually through image augmentation), and aiming to maximize the distance between embeddings of positive pairs and minimize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair for each $x$ in a minibatch $N$ by augmenting each image. Each image in the batch therefore has $12N-$ 1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
@@
    $\ell_{i,j}=\log\sum^{2N}_{k}\exp(sim(z_{i},z_{k}))-sim(z_{i},z_{j})$ (30)
    If we take this loss over an unaugmented batch, i.e the only application of this loss is for $z_{i}$ with itself, then it reduces to:
    $\ell_{i}=\log\sum^{N}_{K}\exp(sim(z_{i},z_{k}))-1$ (31)
    In our representations, the embedding vectors $z$ are one-hot matrices representing a string of characters. For one hot matrices $A_{C\times N}$ and $B_{C\times N}$, the dot product $AB^{T}$ is equivalent to counting the amount of mathcing rows, i.e, the un-normalized Hamming distance, and both matrices have norm $N$:
    $sim(A,B)=A^{T}B/\|A\|\|B\|=\frac{1}{N^{2}}\sum_{i}^{N}\mathds{1}[argmax(A_{i})=argmax(B_{i})]=\frac{C}{N^{2}}(1-d_{H}(A,B))$ (32)
    , in which as before $d_{H}$ is the Hamming distance. We can replace this expresion in the unaugmented NT-Xtent loss:
    $\ell_{i}$ $=$ $\log\sum_{k}^{N}\exp(\frac{C}{N^{2}}(1-d_{H}(z_{i},z_{k}))-1=\log\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))+\frac{C}{N^{2}}-1$
    $\leq$ $\log\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))=\log N+\log\frac{1}{N}\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))$
    $\stackrel{{\scriptstyle\text{(Jensen ineq.)}}}{{\leq}}$ $\log N-\frac{C}{N^{3}}\sum_{k}^{N}d_{H}(z_{i},z_{k}).$
    As a consequence, if $C\leq N^{2}$ then $\ell_{i}\leq\log N-\frac{1}{N}\sum_{k}^{N}d_{H}(z_{i},z_{k})$. Thus, for each element $z_{i}$ in a batch $N$, the negative Hamming distance between $z_{i}$ and all other elements plus the logarithm of the batch size form an upper bound on the NT-Xtent loss for the same element, and minimizing the negative Hamming distance over all elements (or equivalently maximising the positive distance) also minimizes the NT-Xtent loss.
Appendix C Justification of our imitation step starting from classical probabilistic reconstruction games
@@
C.1 Classical probabilistic reconstruction games
Recall that in reconstruction games, the two agent types are a teacher and student (or speaker and listener, or sender and receiver), parameterized by respectively $\theta,\phi$. The game consists in the teacher observing a signal $x$ distributed according to a probability distribution $p$, from which it produces a randomized message $m$ whose distribution conditional on $x$ is obtained according to policy $\pi_{\theta}(\cdot|x)$. The student or listener, observes message $m$ and produces a reconstructed signal distributed according to a conditional probability $\rho_{\phi}(\cdot|m)$. The game dynamics can be modelled by stipulating that together, teacher and student minimize the log-likelihood loss function
@@ C.1 Classical probabilistic reconstruction games
C.2 Reconstruction games in generative tasks
While the above formulation (33) is formalized using the log-likelihood between distributions, for our generative task, the reconstruction modelled slightly differently: the task is to optimize, with high probability, the reconstruction error, measured by a distance between the signal $x$ and its reconstruction $x^{\prime}$, denoted $d(x,x^{\prime})$. Here the notation for $x,x^{\prime}$ and $d(x,x^{\prime})$ can be (ab)used to represent two alternative settings: either (a) $x,x^{\prime}$ are vectors in the autoencoder latent space and $d(x,x^{\prime})$ is a discrepancy (e.g. $d(x,x^{\prime})=|x-x^{\prime}|^{p}$) between them or (b) $x,x^{\prime}$ are reconstructed images and $d(x,x^{\prime})$ is a discrepancy (e.g. based on some ad-hoc norm or otherwise) between them. Thus we have the probabilistic error function

### v3

@@ F.1 Definitions of Holistic and Compositional encodings, and role of the compressibility condition
The compressibility condition on the sender-receiver protocol just requires that the protocol be not holistic, and thus it is fully specified via definitions F.1 and F.2. Furthermore, Definition F.2 fully specifies in which way the map $\mathcal{G}\to\mathcal{M}$ successfully approximates an isomorphic reconstruction of $\mathcal{G}$, if it satisfies conditions compressibility and expressivity.
F.2 Encodings optimizing for only expressivity and efficiency have Kolmogorov Complexity much higher than compositional ones
@@
As a manageable measure of complexity we will use Kolmogorov Complexity (KC), and we verify that in our setting, encodings satisfying expressivity and efficiency will necessarily have higher KC than compositional ones. In particular, the receiver will have more trouble reconstructing from such encodings, justifying our push to enforce the compressibility condition.
Encodings of $\mathcal{D}_{train}$ satisfying expressivity and efficiency include all injective maps
@@
On the other hand, to specify a compositional encoding $\mathsf{comp}$ as in Definition F.2 requires only assigning a string of length $C_{i}$ in alphabet $V$ for every generating factor $G_{i}$, where we can take $C_{i}=\lceil\log_{|V|}N_{i}\rceil$. Thus such string requires $\approx\log_{2}N_{i}$ bits, and the whole encoding has KC given by
@@
Proposition F.3.
Under the same hypotheses on $\mathcal{G}$ as in the previous section, consider the limit $|\mathcal{D}|\to\infty$ and assume that $p=|\mathcal{D}_{train}|/|\mathcal{D}|$ stays bounded away from zero in the limit. Then the expected Kolmogorov complexity of encodings satisfying expressivity and efficiency becomes overwhelmingly larger than the one of compositional encodings:
@@
F.3 Encodings that optimize efficiency under expressivity over $\mathcal{D}_{train}$ are holistic
@@
  • The requirement of "expressivity over $\mathcal{D}_{train}$" means that the encodings are injective over $\mathcal{D}_{train}$.
  • The requirement of "efficiency over $\mathcal{D}_{train}$" means that the encoding maps $\mathcal{D}_{train}$ into $V^{C_{\star}}$ for the minimum value of $C_{\star}$.
@@
For compositional encodings of $\mathbf{G}=[G_{1},\dots,G_{n}]$ the encoding will require to set apart separate message sub-strings of length $C_{i}=\lceil\log_{|V|}N_{i}\rceil$ for factor $G_{i}$ (see Def. F.2), thus the required length for the encoding will be
@@
We now discuss when the inequality signs in (24) are sharp or not, since sharp inequality $C_{0}>C_{\star}$ implies our claim that optimizing for efficiency under expressivity implies non-compositional (i.e., holistic) encodings:
@@
Proposition F.4.
Under the same running assumptions over $\mathcal{G}$, $p$, we have that optimizing only for efficiency of encodings efficiency under injectivity constraints expressivity over $\mathcal{D}_{train}$ is guaranteed to produce holistic (i.e., non-compositional) encodings if one or both of the conditions (25) and (26) holds.
F.4 Possible choices of encoded message regularization
As seen above, if we just optimize for efficiency and expressivity then we are likely to get holistic encodings (Prop. F.4), which will then make it hard for the receiver to generalize the decoding strategy outside the training set due to having overwhelmingly higher expected complexity (Prop. F.3).
This proves that it is important to incentivize condition compressibility, i.e., to push encoding strategies away from overly efficient holistic encodings. In this section we discuss a few approaches to do this in practice, explaining our choice of compressibility regularization.
@@ F.4 Possible choices of encoded message regularization
    Entropy maximization.
    Background and standard approach. Introduced by [57] for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
@@
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in [42].
@@
    Importantly, in our setting, to avoid high variance in estimation due to the large message space (see §G), we approximate the probability distribution of the RNN-produced message token at position $t\leq C$, denoted here $p(m_{t}|x,m_{[t-1]})$, as a $\delta$ distribution over the vocabulary $V$ by using the Gumbel-softmax trick:
@@
    In particular, it becomes clear that optimizing this $h$ becomes unfeasible with gradient methods.
    Our proposal for regularization. Our model can be seen as a different approximation of the entropy of the message. Instead of using the policy weights as probability values, we use the law of large numbers to estimate the probability of each message. For a batch $\mathcal{B}\subset\mathcal{X}$ consisting of $N_{batch}$ training examples and for given sender encoding protocol $S_{\phi}$:
@@
    Yet again, due to the exponential size of the message space, utilizing this probability to estimate the entropy is still not practical. We will assume approximate independence between symbol positions, which allows to get the stronger approximation:
@@
    $H(Pr(m\mid x,\phi))$ =(37)${{=}}-\sum_{t=1}^{C}\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\ln Pr(m_{t}=v\mid x,\phi)$ )
[…]
    =$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)\right.$
    $\quad\left.-\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)^{2}\right)$
    $=\sum_{t=1}^{C}\sum_{v\neq v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)$
    $=\sum_{t=1}^{C}Pr(m_{t}\neq m_{t}^{\prime})=\mathbb{E}[d_{H}(m,m^{\prime})]$
    (36)${{\approx}}\mathbb{E}_{x,x^{\prime}\in\mathcal{B}}[d_{H}(S_{\phi}(x),S_{\phi}(x^{\prime}))],$ (38)
@@
    $\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)$ $=$ $1=\left(\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\right)^{2}$
    $=$ $\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi).$
    The consequence of bound (1) is that maximizing the pairwise Hamming distance between predicted messages in a batch also maximizes a lower bound to the approximate entropy of the message space, justifying the choice of optimizing this quantity (which is easier to i include in practice) for favoring compressibility.
@@
    Contrastive Learning losses. In our framework, PDM can be seen as a form of contrastive loss, which we now show how to connect to standard contranstive losses from previous works.
    Contrastive loss functions such as the triplet loss [52] and NT-Xtent [8] work by defining positive and negative image pairs (usually through image augmentation), and aiming to minimize the distance between embeddings of positive pairs and maximize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair of samples for each $x$ in a mini-batch $N$ by augmenting each image. Each image in the batch therefore has $12N-$ 1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
@@
    $\ell_{i,j}=-sim(z_{i},z_{j})+\log\sum^{2N}_{k}\exp(sim(z_{i},z_{k})).$ (40)
    Here, we apply the same ideas but take this loss over an unaugmented batch, i.e the only application of this loss is for $z_{i}$ against itself. In this case it reduces to:
    $\ell_{i}=-1+\log\sum^{N}_{k}\exp(sim(z_{i},z_{k}))$ (41)
    In our representations, the embedding vectors from a batch $z_{i},i=1,\dots,N$ are $V\times C$ matrices whose columns are approximately one-hot vectors that represent the tokens of a message. For one-hot matrices $z_{i},z_{k}$ of this form, the dot product $z_{i}^{T}z_{k}$ is equivalent to counting the amount of matching rows of $z_{i},z_{k}$, i.e, the un-normalized Hamming distance of the associated message strings $m_{i},m_{k}\in V^{C}$. Since strings have length $C$ and $\|z_{i}\|^{2}=z_{i}^{T}z_{i}\simeq C$, we thus get:
    $sim(z_{i},z_{k})$ $=$ $tr(z_{i}^{T}z_{k})/\|z_{i}\|\|z_{k}\|\simeq\frac{1}{C}\sum_{l=1}^{C}\mathds{1}[\arg\max_{j}((z_{i})_{l,j})=\arg\max_{j}((z_{k})_{l,j})]$
    $\simeq$ $1-d_{H}(m_{i},m_{k}),$
    in which $d_{H}$ is the (normalized by $1/C$) Hamming distance. We can replace this expression in the unaugmented NT-Xtent loss:
    $\ell_{i}$ $=$ $-1+\log\sum_{k=1}^{N}\exp(1-d_{H}(m_{i},m_{k}))$
    $=$ $-1+N-\log\sum_{k=1}^{N}\exp(d_{H}(m_{i},m_{k}))$
    (Jensen $\overset{\begin{subarray}{c}\text{(Jensen}\\ \text{ inequality)}\end{subarray}}{\leq}$ $-1+N-\log N-\frac{1}{N}\sum_{k=1}^{N}d_{H}(m_{i},m_{k}).$
    Then, taking the mean loss over the full batch we get:
    $\frac{1}{N}\sum^{N}_{i=1}\ell_{i}\leq-1+N-\log N-\frac{1}{N^{2}}\sum^{N}_{i=1}\sum^{N}_{k=1}d_{H}(m_{i},m_{k}).$ (42)
    We then include in the loss only the last term, i.e. the average of $d_{H}(m_{i},m_{k})$, which is the only term that actually depends on the batch elements, which forms an approximate upper bound on the NT-Xtent loss.
    Note that our approximation of $sim(z_{i},z_{k})$ by $1-d_{H}(m_{i},m_{k})$, on which the above is based, becomes sharp when the $z_{i}$ have actual one-hot entries, thus the approximation becomes better as our model has high accuracy. Furthermore, Jensen’s inequality, responsbile for the inequality between the term $\ell_{i}$ from NT-Xtent and our metric, becomes closer to equality messages $m_{1},\dots,m_{N}$ become uniformly spread across $V^{C}$, which is the case for minimizers of our metric as well. This means that, at least heuristically speaking, we can consider our metric as having similar minimizers to the NT-Xtent regularization.
    A precise matching lower bound to (42) is left for future work.
Appendix G Justification of our imitation step starting from classical probabilistic reconstruction games
@@
G.1 Classical probabilistic reconstruction games
Recall that in reconstruction games, the two agent types are a teacher and student (or speaker and listener, or sender and receiver), parameterized by respectively $\theta,\phi$. The game consists in the teacher observing a signal $x$ distributed according to a probability distribution $p$, from which it produces a randomized message $m$ whose distribution conditional on $x$ is obtained according to policy $\pi_{\theta}(\cdot|x)$. The student or listener, observes message $m$ and produces a reconstructed signal distributed according to a conditional probability $\rho_{\phi}(\cdot|m)$. The game dynamics can be modeled by stipulating that together, teacher and student minimize the log-likelihood loss function
@@ G.1 Classical probabilistic reconstruction games
G.2 Reconstruction games in generative tasks
While the above formulation (43) is formalized using the log-likelihood between distributions, for our generative task, the reconstruction modeled slightly differently: the task is to optimize, with high probability, the reconstruction error, measured by a distance between the signal $x$ and its reconstruction $x^{\prime}$, denoted $d(x,x^{\prime})$. Here the notation for $x,x^{\prime}$ and $d(x,x^{\prime})$ can be (ab)used to represent two alternative settings: either (a) $x,x^{\prime}$ are vectors in the autoencoder latent space and $d(x,x^{\prime})$ is a discrepancy (e.g. $d(x,x^{\prime})=|x-x^{\prime}|^{p}$) between them or (b) $x,x^{\prime}$ are reconstructed images and $d(x,x^{\prime})$ is a discrepancy (e.g. based on some ad-hoc norm or otherwise) between them. Thus we have the probabilistic error function